A company is considering launching a new product, this product is called X. This product X includes a material called Alpha. This material is also included in today's main product Y. The following data have been produced as the basis for the decision:

Direct material for Y: £40/piece
Direct pay for Y: 18/piece
Overhead charge for Y: 45/piece
Self-cost (sum) for: 103/piece

Direct material for X: £250/piece
Direct pay for X: 350/piece
Overhead charge for X: 630/piece
Self-cost (sum) for X: 1230/piece

Of the direct material cost of Y, the Alpha is responsible for £10. The corresponding cost for X is £60. The sale price of Y is £120 and the company is currently selling everything you can manufacture. For the foreseeable future, it is not possible to get more of the Alpha material.

What is the lowest acceptable price for X under these conditions?
Answer #1

Total cost for Y= 103 (of which 10 is alpha cost)

Selling rate for Y= 120.

Profit for 1 piece sale of Y= 120-103= 17.

ie, the company makes 17 profit for each 10 worth of Alpha consumed. or 1.7/per unit alpha.

Assuming that Alpha is the only exhaustible resource and all other resources are non depleting, Alpha is the bottle neck for company profits. Hence, Alpha should produce more than 1.7 profit per unit consumption, when used in product X. A single piece of X consumes alpha worth 60. Hence, the profit generated by that unit should atleast be worth (1.7 x 60)= 102.

If the cost is 1230, the lowest acceptable price would be: 1230+102= £1332